# Global identification of dynamic panel models with interactive effects

### v1

@@ 3 Assumptions and useful results
Assumption 1.
The number of distinct factors is known and equal to $\bar{r}$
Assumption 2.
Identification restrictions IC1 or IC2:
IC1: $\mathbf{\Psi}$ an unrestricted PD matrix and $\mathbf{F}=(\mathbf{I}_{\bar{r}},\mathbf{F}_{2}^{{}^{\prime}})^{\prime}$
IC2: $\mathbf{\Psi}$ is diagonal and $T^{-1}\mathbf{F}^{{}^{\prime}}\mathbf{D^{-1}}\mathbf{F}=\mathbf{I}_{\bar{r}}$
This assumption comes from the fact that factor models (and thus interactive effects) are identifiable up to a $\bar{r}\times\bar{r}$ matrix rotation, so $\bar{r}^{2}$ restrictions must be imposed for identification. A discussion of why this is the case can be found in Section 5 of Anderson and Rubin (1956). We will work with IC1 for the rest of the paper. But it can be shown that the model is globally identifiable under IC2 (with some modifications).
Assumption 3.
$T\geq 2(\bar{r}+1)$
Assumption 4.
$\lambda_{i}$ are either random or fixed constants with $\bm{\Psi}_{N}\to\bm{\Psi}$ as $N\to\infty$.
Assumption 5.
$\alpha,\mathbf{\Psi},d_{1},d_{2},\dots,d_{T}$ are fixed population parameters.
Assumption 6.
Let $\mathbf{F}_{2}=\begin{bmatrix}f_{\bar{r}+1,1}&f_{\bar{r}+1,2}&\dots&f_{\bar{r}+1,\bar{r}}\\ \vdots&\vdots&\vdots&\vdots\\ f_{T,1}&f_{T,2}&\dots&f_{T,\bar{r}}\end{bmatrix}$ be the random elements of $\mathbf{F}$. Then, $\mathbf{F}_{2}$ is a continuous random matrix with support in $\Xi\in\mathbb{R}^{(T-\bar{r})\times\bar{r}}$.
Additionally, throughout the paper, we will reference several established results from linear algebra and probability theory. To ensure they are easily accessible to the reader, we will present them as lemmas in this section.
@@
where $\mathbf{A}_{-i,-j}$ is a minor of $\mathbf{A}$ obtained from removing row $i$ and column $j$ of matrix $\mathbf{A}$, and $A_{i,j}$ represents the $(i,j)$-th element of the matrix $\mathbf{A}$.
In this paper, we will apply Lemma 3 for cases in which $\mathbf{A}$ itself is a minor of a larger matrix. Let $\mathbf{H}$ be a $T\times T$ matrix, and let $\mathbf{M}^{\mathbf{H}}_{R,C}$ be a minor of $\mathbf{H}$ of dimension $k$ including rows $R=\{r_{1},\dots,r_{k}\}$ and columns $C=\{c_{1},\dots,c_{k}\}$ of the original matrix. Therefore, for any $r\in R$ the Laplace decomposition of this minor is given by:
$\det(\mathbf{M}^{\mathbf{H}}_{R,C})=\sum_{j=1}^{k}(-1)^{r+j}H_{r,c_{j}}\det\left(\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}\right)$
where $H_{r,c_{j}}$ is the $(r,c_{j})$-th element of $\mathbf{H}$, and $\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}$ is the minor [^2] Where $R-r$ means removal $r$ from the set $R$, and $C-c_{j}$ means removal of $c_{j}$ from the set $C$. They correspond to the usual notation $R\setminus\{r\}$ and $C\setminus\{c_{j}\}$. of $\mathbf{H}$ that includes rows $R-r$, and columns $C-c_{j}$.
@@ 4 Global identification
In this section, we demonstrate that, under Assumptions 1-6, the parameters of the dynamic panel data model with interactive effects described in (1) are globally identified.
@@ 4.1 Proof insight
The proof we propose draws inspiration from Theorem 5.1 of Anderson and Rubin (1956). In that theorem, the authors establish a sufficient condition for global identification of a factor model that did not include an autoregressive component. In their simpler model, global identification requires
@@
Then, they leverage the fact that all determinants of the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$ must be zero for (6) to hold (Lemma 1). This follows as a necessary condition because, by assumption, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is of rank $\bar{r}$. This procedure enables them to pin-down the diagonal terms of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$, thereby establishing that $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$. Once this is established, it follows immediately that $\mathbf{\widetilde{D}}=\mathbf{D}$, completing the proof.
@@
Once we establish that $\alpha$ is globally identified, it follows that our identification condition can be rewritten in the form proposed by Anderson and Rubin (1956), ensuring that all other parameters can also be globally identified.
Thus, the primary focus of our analysis is to show that $\alpha$ is globally identified (or identically, that it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to hold). This task requires some intermediate results, which we will articulate and prove throughout the paper. The main contribution of this work is to formally establish the following Theorem.
@@
Let $y_{it}$ be generated from (1). Then, under Assumptions 1-6, $\alpha$ is almost surely globally identified.
Once this shown, it is immediate to conclude that all the parameters of the model are almost surely globally identified as well, which is formally stated in the following Theorem
@@
Let $y_{it}$ be generated from (1). Then, under Assumptions 1-6, all the parameters in the model are almost surely globally identified.
Remark: The almost sure identification refers to the distribution of $\mathbf{F}$, not to $\alpha$. Any value of $\alpha$ is globally identifiable. The practical implication is that as long as there is sufficient variation in $f_{t}$, both $\alpha$ and other parameters are globally identifiable. This is particularly true when $T$ is sufficiently large, ensuring a sufficient number of distinct realizations of $f_{t}$.
@@ 4.2 Intermediate results
In order to prove Theorem 1, we will rearrange the problem in a way that resembles the proof of Anderson and Rubin (1956) and then establish some properties of this modified problem. Start from (5). As $\mathbf{\widetilde{\Gamma}}$ is a unit-diagonal lower triangular matrix, it is invertible so that we can pre-multiply and post-multiply by $\mathbf{\widetilde{\Gamma}^{-1}}$ and $\mathbf{\widetilde{\Gamma}^{-1^{\prime}}}$ and re-write
@@
where
@@
Proof: We have shown that (5) is equivalent to $\mathbf{O}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$. Furthermore, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD matrix of rank $\bar{r}$ by assumption. Then, it must be the case that $\mathbf{O}$ is also a symmetric PSD matrix of rank $\bar{r}$ in order for (5) to hold. By Lemma 1 we know that if $\mathbf{O}$ is of rank $\bar{r}$ then the determinant of all the minors of $\mathbf{O}$ of dimension $k>\bar{r}$ must be equal to 0. $\square$
@@
$\mathbf{O}$ can be decomposed into $\mathbf{O}=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi}),$
where $\mathbf{\Omega}$ is defined as in (3) and $\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi})$ is given by:
@@
Proof: By definition $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F^{\prime}}+\mathbf{D}$, so we can write $\mathbf{O}$ as
@@
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\alpha,\widetilde{\alpha},\mathbf{D},\mathbf{F},\mathbf{\Psi})$
where $\mathbf{J}(\alpha,\widetilde{\alpha},\mathbf{D},\mathbf{F},\mathbf{\Psi})=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ This proves the claim. $\Box$
From this point forward, we will use either $\mathbf{J}$ or $\mathbf{J}\left(\widetilde{\alpha},\theta\right)$ to refer to $\mathbf{J}(\widetilde{\alpha},\alpha,\mathbf{D},\mathbf{F},\mathbf{\Psi})$, noting that it depends on the true data generating process parameters $\theta=\text{vec}(\alpha,\mathbf{F},\mathbf{\Psi},\mathbf{D})$ and on $\widetilde{\alpha}$. With this result in hand, we can proceed to derive the following corollaries.
@@
Proof: As $\mathbf{\widetilde{D}}$ is a diagonal matrix then $\widetilde{D}_{r,c}=0$ for all $r\neq c$. Using the expression for $\mathbf{O}$ derived in Claim 2:
@@
For every row index $r\in\{\bar{r}+1,\dots T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $c\neq 1$. Similarly for every column index $c\in\{\bar{r}+1,\dots,T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $r\neq 1$.
Proof: We only provide the proof for the first statement, as the second one is completely analogous given the symmetry of the problem.
To show this result, we will use the fact that most of the positions of $\mathbf{\Omega}$ are non-zero with probability 1. First, note that under IC1 (Assumption 2) we know that the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ is given by:
@@
where $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ represents the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ (same for $\mathbf{D}_{1:\bar{r},1:\bar{r}}$). Therefore, we cannot guarantee that the positions of $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ are non-zero, as the covariance between the factor loadings (the off-diagonal elements of $\mathbf{\Psi}$) could be equal to zero. Then, we will use information contained in the remaining elements of $\mathbf{\Omega}$ to show this result.
@@
From the definition of $\mathbf{\Omega}$ under IC1 and Assumption 6 we know that $\Omega_{r,c}\neq 0$ with probability 1 if $r>\bar{r}$, as every position depends on at least one factor realization, so the case in which the position is exactly equal to zero is an event of probability zero (Lemma 2). Now, recall that $\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ is a $T\times T$ matrix. By Assumptions 5 and 6 we know that the elements of $\mathbf{\Omega}_{(\bar{r}+1):T,1:T}$ follow a continuous distribution dependent on $\mathbf{F}$. We establish the corollary in three steps.
@@
First, we show that $\mathbf{\Omega}\mathbf{L}^{\prime}$ is a random matrix with non-zero elements for rows $r>\bar{r}$ and $c\neq 1$. To see this, observe that
$\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}\Omega_{r,i}L^{{}^{\prime}}_{i,c}=\sum_{i=1}^{T}\Omega_{r,i}L_{c,i}$
@@
If $c=1$ then we know that $L_{1,i}=0$ for all $i$, implying that $\mathbf{\Omega}\mathbf{L}^{\prime}_{r,1}=0$. Conversely, for $c>1$, the elements $L_{c,c-1},L_{c,c-2},\dots L_{c,1}$ are all nonzero if $\alpha\neq 0$. If $\alpha=0$ we still have that $L_{c,c-1}=1$, which suffices for our purpose (the element $\Omega_{r,c-1}$ is non-zero with probability 1 when $r>\bar{r}$). Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indexes $r>\bar{r}$, it follows that $\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}\neq 0$ with probability 1 if $c\neq 1$.
@@
By the same reasoning, the matrix $\mathbf{L}\mathbf{\Omega}$ has zeros in the first row and nonzero entries $\mathbf{L}\mathbf{\Omega}_{r,c}$ for every column $c>\bar{r}$ with probability 1 if $r\neq 1$.
@@
Third, we show that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has non-zero elements for rows $r>\bar{r}+1$ and $c\neq 1$ with probability 1. Using matrix associativity,
$\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}L_{r,i}(\Omega L^{{}^{\prime}})_{i,c}$
@@
Since $L_{1,i}=0$ for all $i$ and $(\Omega L^{{}^{\prime}})_{i,1}=0$ for all $i$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has zeros in the first row and first column. Moreover, for row indexes $r>\bar{r}$ we know that the elements $L_{r,r-1},L_{r,r-2},\dots L_{r,1}$ are nonzero if $\alpha\neq 0$. Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indexes $r>\bar{r}$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{{}^{\prime}}_{r,c}\neq 0$ with probability 1 if $\alpha\neq 0$. If $\alpha=0$ then we require $r>\bar{r}+1$ as $L_{r,r-2},\dots L_{r,1}$ are all equal to zero, so only $L_{r,r-1}=1$, but if $r=\bar{r}+1$ then $(\Omega L^{{}^{\prime}})_{r-1,c}=(\Omega L^{{}^{\prime}})_{\bar{r},c}$ may be equal to zero (the results derived before for $\mathbf{\Omega}\mathbf{L}^{\prime}$ hold under $r>\bar{r}$). Then, in the most general case, if $r>\bar{r}+1$ and $c\neq 1$ then $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1.
@@
Recall that $J_{r,c}=(\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}+(\mathbf{L}\mathbf{\Omega})_{r,c}+(\alpha-\widetilde{\alpha})(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$. We have shown that the first element is non-zero with probability 1 if $r>\bar{r}$, the second element is non-zero with probability 1 if $c>\bar{r}$ and $(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$ is non-zero with probability 1 if $r>\bar{r}+1$. It is clear that in any case the event in which the summands cancel out can only occur in a set of factor realizations that occur with probability zero (Lemma 2 and Assumption 6). This completes the proof. $\square$
@@
The minors in the definition are squared matrices, thus $|R|=|C|=k$. They do not contain diagonal element, thus $R\cap C=\emptyset$. It follows that $|R|+|C|=2k\leq T$.

### v2

@@ 3 Assumptions and useful results
Assumption A1.
The number of distinct factors is known and equal to $\bar{r}$.
Assumption A2.
$\mathbf{\Psi}$ is an unrestricted symmetric and PD matrix, and $\mathbf{F}=(\mathbf{I}_{\bar{r}},\mathbf{F}_{2}^{\prime})^{\prime}$.
This assumption comes from the fact that factor models (and thus interactive effects) are identifiable up to a $\bar{r}\times\bar{r}$ matrix rotation, so $\bar{r}^{2}$ restrictions must be imposed for identification. A discussion of why this is the case can be found in Section 5 of Anderson and Rubin, (1956). We work with this restriction for convenience, but any other normalization that imposes $\bar{r}^{2}$ restriction would also work.
Assumption A3.
$T\geq 2(\bar{r}+1)$.
Assumption A4.
$\varepsilon_{i}$ are i.i.d. over $i$. $\mathbb{E}[\varepsilon_{it}]=0$, $\mathbb{V}(\varepsilon_{it})=d_{t}>0$, and $\varepsilon_{it}$ are independent over $t$. $\varepsilon_{it}$ is independent from the factor loadings $\lambda_{i}$ and the factors $\mathbf{F}$.
Assumption A5.
$\lambda_{i}$ are either random or fixed constants with $\mathbf{\Psi}_{N}\to\mathbf{\Psi}$ as $N\to\infty$.
Assumption A6.
$\alpha,d_{1},d_{2},\dots,d_{T},\mathbf{\Psi}$ are fixed population parameters.
Assumption A7.
Let $\mathbf{F}_{2}=\begin{bmatrix}f_{\bar{r}+1,1}&f_{\bar{r}+1,2}&\dots&f_{\bar{r}+1,\bar{r}}\\ \vdots&\vdots&\vdots&\vdots\\ f_{T,1}&f_{T,2}&\dots&f_{T,\bar{r}}\end{bmatrix}$ be the random elements of $\mathbf{F}$. Then, $\mathbf{F}_{2}$ is a continuous random matrix with support in $\Pi\in\mathbb{R}^{(T-\bar{r})\times\bar{r}}$.
Additionally, throughout the paper, we will reference several established results from linear algebra and probability theory. To ensure they are easily accessible to the reader, we present them as lemmas in this section.
@@
where $\mathbf{A}_{-i,-j}$ is the minor obtained by removing row $i$ and column $j$, and $A_{i,j}$ represents the $(i,j)$-th element of the matrix $\mathbf{A}$.
In this paper, we will apply Lemma 3 for cases in which $\mathbf{A}$ itself is a minor of a larger matrix. Let $\mathbf{H}$ be a $T\times T$ matrix, and let $\mathbf{M}^{\mathbf{H}}_{R,C}$ be a minor of $\mathbf{H}$ of dimension $k$ including rows $R=\{r_{1},\dots,r_{k}\}$ and columns $C=\{c_{1},\dots,c_{k}\}$ of the original matrix. Therefore, for any $r\in R$ we can denote $i\in\{1,\dots,k\}$ as the position of row $r$ in the minor $\mathbf{M}^{\mathbf{H}}_{R,C}$, and the Laplace decomposition of this minor is given by:
$\det(\mathbf{M}^{\mathbf{H}}_{R,C})=\sum_{j=1}^{k}(-1)^{i+j}H_{r,c_{j}}\det\left(\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}\right)$
where $H_{r,c_{j}}$ is the $(r,c_{j})$-th element of $\mathbf{H}$, and $\mathbf{M}^{\mathbf{H}}_{R-r,C-c_{j}}$ is the minor [^3] Where RrR-r means removal $r$ from the set $R$, and $C-c_{j}$ means removal of $c_{j}$ from the set $C$. They correspond to the usual notation $R\setminus\{r\}$ and $C\setminus\{c_{j}\}$. of $\mathbf{H}$ that includes rows RrR-r, and columns $C-c_{j}$.
@@ 4 Global identification
In this section, we demonstrate that, under Assumptions A1-A7, the parameters of the dynamic panel data model with interactive effects described in (1) are globally identified.
@@ 4.1 Proof insight
The proof we propose draws inspiration from Theorem 5.1 of Anderson and Rubin, (1956). In that theorem, the authors establish a sufficient condition for global identification of a factor model that did not include an autoregressive component. In their simpler model, global identification requires
@@
Then, they leverage the fact that all determinants of the $(\bar{r}+1)\times(\bar{r}+1)$ minors of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}+\mathbf{\widetilde{D}}-\mathbf{D}$ must be zero for (6) to hold (Lemma 1). This follows as a necessary condition because, by assumption, $\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$ is of rank $\bar{r}$. This procedure enables them to pin down the diagonal terms of $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$, thereby establishing that $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}$. Once this is established, it follows immediately that $\mathbf{\widetilde{D}}=\mathbf{D}$, completing the proof.
@@
Once we establish that $\alpha$ is globally identified, it follows that our identification condition can be rewritten in the form proposed by Anderson and Rubin, (1956), ensuring that all other parameters can also be globally identified.
Thus, the primary focus of our analysis is to show that $\alpha$ is globally identified (or equivalently, that it must be the case that $\widetilde{\alpha}=\alpha$ in order for (5) to hold). This task requires some intermediate results, which we will articulate and prove throughout the paper. The main contribution of this work is to formally establish the following Theorem.
@@
Let $y_{it}$ be generated from (1). Then, under Assumptions A1-A7, $\alpha$ is almost surely globally identified.
Once this is shown, it is immediate to conclude that all the parameters of the model are almost surely globally identified as well, which is formally stated in the following Theorem
@@
Let $y_{it}$ be generated from (1). Then, under Assumptions A1-A7, all the parameters in the model are almost surely globally identified.
Remark: The qualifier ‘almost surely’ is with respect to the distribution of $\mathbf{F}$. For any fixed $\alpha$, identification holds for almost all realizations of $\mathbf{F}$. The practical implication is that as long as there is sufficient variation in $f_{t}$, both $\alpha$ and other parameters are globally identifiable. This is particularly true when $T$ is sufficiently large, ensuring a sufficient number of distinct realizations of $f_{t}$.
@@ 4.2 Intermediate results
In order to prove Theorem 1, we will rearrange the problem in a way that resembles the proof of Anderson and Rubin, (1956) and then establish some properties of this modified problem. Start from (5). As $\mathbf{\widetilde{\Gamma}}$ is a unit-diagonal lower triangular matrix, it is invertible so that we can pre-multiply and post-multiply by $\mathbf{\widetilde{\Gamma}}^{-1}$ and $\mathbf{\widetilde{\Gamma}}^{-1^{\prime}}$ and re-write
@@
where $\mathbf{L}$ is a strictly lower-triangular Toeplitz matrix given by
@@
Proof.
We have shown that (5) is equivalent to $\mathbf{O}=\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$. Furthermore, $\mathbf{\widetilde{F}}\mathbf{\widetilde{\Psi}}\mathbf{\widetilde{F}}^{\prime}$ is a symmetric PSD matrix of rank $\bar{r}$ by assumption. Then, it must be the case that $\mathbf{O}$ is also a symmetric PSD matrix of rank $\bar{r}$ in order for (5) to hold. By Lemma 1 we know that if $\mathbf{O}$ is of rank $\bar{r}$ then the determinant of all the minors of $\mathbf{O}$ of dimension $k>\bar{r}$ must be equal to 0. $\square$
@@
$\mathbf{O}$ can be decomposed into $\mathbf{O}=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\theta^{0}),$
where $\mathbf{\Omega}$ is defined as in (3) and $\mathbf{J}(\widetilde{\alpha},\theta^{0})$ is given by:
@@
Proof.
By definition $\mathbf{\Omega}=\mathbf{F}\mathbf{\Psi}\mathbf{F}^{\prime}+\mathbf{D}$, so we can write $\mathbf{O}$ as
@@
$=\mathbf{\Omega}-\mathbf{\widetilde{D}}+(\alpha-\widetilde{\alpha})\mathbf{J}(\widetilde{\alpha},\theta^{0})$
where $\mathbf{J}(\widetilde{\alpha},\theta^{0})=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}.$ This proves the claim. $\square$
With this result in hand, we can proceed to derive the following corollaries.
@@
Proof.
As $\mathbf{\widetilde{D}}$ is a diagonal matrix then $\widetilde{D}_{r,c}=0$ for all $r\neq c$. Using the expression for $\mathbf{O}$ derived in Claim 2:
@@
For every row index $r\in\{\bar{r}+1,\dots T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $c\geq 2$. Similarly for every column index $c\in\{\bar{r}+1,\dots,T\}$ we have that $J_{r,c}\neq 0$ with probability 1 if $r\geq 2$.
Proof.
We only provide the proof for the first statement, as the second one is completely analogous given the symmetry of the problem.
To show this result, we use the fact that most of the positions of $\mathbf{\Omega}$ are non-zero with probability 1. First, note that by Assumption A2 we know that the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ is given by:
@@
where $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ represents the first $\bar{r}\times\bar{r}$ block of $\mathbf{\Omega}$ (same for $\mathbf{D}_{1:\bar{r},1:\bar{r}}$). Therefore, we cannot guarantee that the positions of $\mathbf{\Omega}_{1:\bar{r},1:\bar{r}}$ are non-zero, as the covariance between the factor loadings (the off-diagonal elements of $\mathbf{\Psi}$) could be equal to zero. Then, we use information contained in the remaining elements of $\mathbf{\Omega}$ to show this result.
@@
From the definition of $\mathbf{\Omega}$ under Assumption A2 and Assumption A7 we know that $\Omega_{r,c}\neq 0$ with probability 1 if $r>\bar{r}$, as every position depends on at least one factor realization, so the case in which the position is exactly equal to zero is an event of probability zero (Lemma 2). Now, recall that $\mathbf{J}=\mathbf{L}\mathbf{\Omega}+\mathbf{\Omega}\mathbf{L}^{\prime}+(\alpha-\widetilde{\alpha})\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ is a $T\times T$ matrix. By Assumptions A6 and A7 we know that the elements of $\mathbf{\Omega}_{(\bar{r}+1):T,1:T}$ follow a continuous distribution dependent on $\mathbf{F}$. We establish the corollary in three steps.
@@
First, we show that $\mathbf{\Omega}\mathbf{L}^{\prime}$ is a random matrix with non-zero elements for rows $r>\bar{r}$ and $c\neq 1$. From the definition of $\mathbf{L}$ in 2 we know that:
$L_{r,c}=\begin{cases}0,\quad r\leq c\\ 1,\quad r=c+1\\ \alpha^{r-c-1},\quad r>c+1\end{cases}$
Then, observe that
$(\mathbf{\Omega}\mathbf{L}^{\prime})_{r,c}=\sum_{i=1}^{T}\Omega_{r,i}L^{\prime}_{i,c}=\sum_{i=1}^{T}\Omega_{r,i}L_{c,i}=\sum_{i=1}^{c-1}\Omega_{r,i}\alpha^{c-i-1}$
@@
where from the definition of $\mathbf{L}$ we know that all indexes $i\geq c$ are such that $L_{c,i}=0$. For $c\geq 2$, the elements $L_{c,c-1},L_{c,c-2},\dots L_{c,1}$ are all nonzero if $\alpha\neq 0$. If $\alpha=0$ we still have that $L_{c,c-1}=1$, which suffices for our purpose (the element $\Omega_{r,c-1}$ is non-zero with probability 1 when $r>\bar{r}$). Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indices $r>\bar{r}$, it follows that $\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1 if $c\geq 2$.
@@
By the same reasoning, the matrix $\mathbf{L}\mathbf{\Omega}$ has zeros in the first row and nonzero entries $\mathbf{L}\mathbf{\Omega}_{r,c}$ for every column $c>\bar{r}$ with probability 1 if $r\geq 2$.
@@
Third, we show that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has non-zero elements for rows $r>\bar{r}+1$ and $c\geq 2$ with probability 1. Using matrix associativity,
$\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}=\sum_{i=1}^{T}L_{r,i}(\Omega L^{\prime})_{i,c}$
@@
Since $L_{1,i}=0$ for all $i$ and $(\Omega L^{\prime})_{i,1}=0$ for all $i$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}$ has zeros in the first row and first column. Moreover, for row indices $r>\bar{r}$ we know that the elements $L_{r,r-1},L_{r,r-2},\dots L_{r,1}$ are nonzero if $\alpha\neq 0$. Then, by Lemma 2 and the definition of $\mathbf{\Omega}$ for row indices $r>\bar{r}$, it follows that $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1 if $\alpha\neq 0$. If $\alpha=0$ then we require $r>\bar{r}+1$ as $L_{r,r-2},\dots L_{r,1}$ are all equal to zero, so only $L_{r,r-1}=1$, but if $r=\bar{r}+1$ then $(\Omega L^{\prime})_{r-1,c}=(\Omega L^{\prime})_{\bar{r},c}$ may be equal to zero (the results derived before for $\mathbf{\Omega}\mathbf{L}^{\prime}$ hold under $r>\bar{r}$). Then, in the most general case, if $r>\bar{r}+1$ and $c\geq 2$ then $\mathbf{L}\mathbf{\Omega}\mathbf{L}^{\prime}_{r,c}\neq 0$ with probability 1.
@@
Recall that $J_{r,c}=(\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}+(\mathbf{L}\mathbf{\Omega})_{r,c}+(\alpha-\widetilde{\alpha})(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$. We have shown that the first element is non-zero with probability 1 if $r>\bar{r}$, the second element is non-zero with probability 1 if $c>\bar{r}$ and $(\mathbf{L}\mathbf{\Omega}\mathbf{L^{\prime}})_{r,c}$ is non-zero with probability 1 if $r>\bar{r}+1$. The event in which the summands cancel out can only occur in a set of factor realizations that occur with probability zero (Lemma 2 and Assumption A7). This completes the proof. $\square$
@@
The minors in the definition are squared matrices, thus $|R|=|C|=k$. They do not contain diagonal elements, thus $R\cap C=\emptyset$. It follows that $|R|=|C|=k\leq\frac{T}{2}$.